What is the PH of a solution whose hydroxide

-3

ion (OH) is 0.003 moldm


1
Expert's answer
2022-05-30T17:05:04-0400

Solution:

[OH] = 0.003 mol dm−3 = 0.003 M


We can convert between [OH] and pOH using the following equation:

pOH = −log[OH] = −log(0.003) = 2.52


For any aqueous solution at 25C:

pH + pOH = 14

pH = 14 − pOH = 14 − 2.52 = 11.48

pH = 11.48


Answer: pH = 11.48